# SpatialTree: How Spatial Abilities Branch Out in MLLMs

### v1

@@ C.2 Agentic Competence (L4)
Spatial Action Mapping. In the context of spatial agents, navigation and manipulation represent the most common forms of interaction within 3D environments. We address each with a distinct action space design. For navigation, we conceptualize agent movement as a series of camera motion controls (referring to recent video world models [4, 35]). To enable precise and intuitive control, we decompose complex camera movements into a set of fundamental motion primitives inspired by established cinematography techniques. This approach allows us to translate high-level language instructions (e.g., "move to the left," "look up") into a structured, low-level action space. The six fundamental primitives, their corresponding cinematic terms, degrees of freedom (DoF), and parameterization are detailed in Tab. A.
@@
Goal-driven Manipulation For the Goal-Driven Manipulation capability, we utilize action annotations from the Droid [22] and EgoDex [16] datasets. This task requires the MLLM to generate a sequence of precise actions to move a robot end-effector or a human hand from a starting state to a target state, both specified by images. The action space for Droid encompasses 7-DoF control: 6-DoF for the end-effector’s pose (translation and rotation) and a binary state for the gripper (open/close). A similar action space is adapted for EgoDex, controlling wrist pose and finger grip. The MLLM is prompted to generate a sequence of continuous action vectors, as shown in the template below:

### v2

@@ C.2 Agentic Competence (L4)
Spatial Action Mapping. In the context of spatial agents, navigation and manipulation represent the most common forms of interaction within 3D environments. We address each with a distinct action space design. For navigation, we conceptualize agent movement as a series of camera motion controls (referring to recent video world models [genie3, yume]). To enable precise and intuitive control, we decompose complex camera movements into a set of fundamental motion primitives inspired by established cinematography techniques. This approach allows us to translate high-level language instructions (e.g., "move to the left," "look up") into a structured, low-level action space. The six fundamental primitives, their corresponding cinematic terms, degrees of freedom (DoF), and parameterization are detailed in Tab. A.
@@
Goal-driven Manipulation For the Goal-Driven Manipulation capability, we utilize action annotations from the Droid [khazatsky2024droid] and EgoDex [egodex] datasets. This task requires the MLLM to generate a sequence of precise actions to move a robot end-effector or a human hand from a starting state to a target state, both specified by images. The action space for Droid encompasses 7-DoF control: 6-DoF for the end-effector’s pose (translation and rotation) and a binary state for the gripper (open/close). A similar action space is adapted for EgoDex, controlling wrist pose and finger grip. The MLLM is prompted to generate a sequence of continuous action vectors, as shown in the template below: